OpenStudy (anonymous):

Which function has the greatest average rate of change on the interval [1, 5]?

myininaya (myininaya):

for each function you have to choose from compute the slope from x=1 to x=5

myininaya (myininaya):

\[\frac{f(5)-f(1)}{5-1}=\frac{f(5)-f(1)}{4}\]

OpenStudy (anonymous):

y = 4x^2?

OpenStudy (anonymous):

* y=4^x* thanks!